Ammonium phosphate NH43PO4 is an important ingredient in many fertilizers. It can be made by reacting phosphoric acid H3PO4 with ammonia NH3. What mass of ammonium phosphate is produced by the reaction of 10.g of ammonia? Be sure your answer has the correct number of significant digits.


1
Expert's answer
2021-03-26T05:37:48-0400

Solution:

The balanced chemical equation:

3NH3 + H3PO4 → (NH4)3PO4

According to the equation above: n(NH3)/3 = n((NH4)3PO4)


Moles of NH3 = Mass of NH3 / Molar mass of NH3

The molar mass of NH3 is 17.031 g mol-1.

Hence,

Moles of NH3 = 10.0 g / 17.031 g mol-1 = 0.5872 mol


n((NH4)3PO4) = n(NH3)/3 = 0.5872 mol / 3 = 0.1957 mol

Mass of (NH4)3PO4 = Moles of (NH4)3PO4 × Molar mass of (NH4)3PO4

The molar mass of (NH4)3PO4 is 149.087 g mol-1.

Hence,

Mass of (NH4)3PO4 = 0.1957 mol × 149.087 g mol-1 = 29.18 g = 29. g

Mass of (NH4)3PO4 = 29 g


OR (short form solution):

(10. g NH3) × (1 mol NH3 / 17.031 g NH3) × (1 mol (NH4)3PO4 / 3 mol NH3) × (149.087 g (NH4)3PO4 / 1 mol (NH4)3PO4) = 29.18 g (NH4)3PO4 = 29. g (NH4)3PO4


Answer: 29. g of ammonium phosphate is produced.